OpenStudy (curiouslocal):

can anyone help me? i'm confused: csc^2Atan^2A - 1 = tan^2A

OpenStudy (mww):

\[\csc^2A = \frac{ 1 }{ \sin^2A }\] \[\tan^2A = \frac{ \sin^2A }{ \cos^2A }\] \[LHS = \csc^2A \tan^2A - 1 = \frac{ 1 }{ \sin^2A } \frac{ \sin^2A }{ \cos^2A } - 1= \frac{ 1 }{ \cos^2A } - 1 = \frac{ 1-\cos^2A }{ \cos^2A }\] \[= \frac{ \sin^2A }{ \cos^2A } = \tan^2A = \] alternatively from 1/cos^2A - 1 can use the Pythagorean trig identity sec^2A - 1 = tan^2A \[\frac{ 1 }{ \cos^2A } - 1 \]

OpenStudy (princeharryyy):

do u have to find A in this?

OpenStudy (curiouslocal):

no.. just proving it

OpenStudy (princeharryyy):

or just prove that this is right? that LHS is equal to RHS

OpenStudy (princeharryyy):